# Algorithms for Optimizing Acyclic Queries

### v1

@@ 2 Related Work
Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [28, 34, 30], cost-based pruning from the top down [11, 15], greedy heuristics [5, 14, 38], and randomized search [36]. Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [21, 23], while others avoid Cartesian products [28, 29]. Our algorithms restrict the query plans to those running in linear time for acyclic queries. In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^1] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other. We later define this as the line graph of the query hypergraph in definition 2. our algorithms find maximum spanning trees of the weighted join graph. Several algorithms for ordering binary joins are based on dynamic programming and tabulate shared structures among different plans [28, 29]. This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner. Our enumeration algorithm avoids redundant work by enumerating join trees by edits, i.e., it outputs the difference between consecutive join trees instead of the trees themselves. Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
@@ 13 $R\leftarrow R\setminus\{r_{i}\}$
fig.˜2(a) shows an example hypergraph of size 14. The line graph $L_{6}=L(H_{6})$ is shown in fig.˜2(b). For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph. The edge weight is given by $\omega(Y,U)=|\{a,d\}|=2$. Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$. Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge. The size of the line graph is $|L_{6}|=19$.
@@
The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree. It generates the join tree $T^{G}$ as shown in fig.˜2(c). Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic. The detailed definitions are given by definition˜41 in appendix˜A.
Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40]. algorithm˜1 shows a simplified pseudocode of MCS. We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in fig.˜2(d) with example˜42 in appendix˜A.
@@ 4.1 Equivalent Graph
Figure 2: $T_{W}$ is an MST of the line graph in fig.˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges. $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
@@
Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in example˜10. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
@@
Given the 6-clique line graph $L_{6}$ in fig.˜2(b), we find a rooted MST $T_{W}$ as shown in fig.˜2(a). Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$. We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in fig.˜2(b). Non-tree edges can slide similarly. We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in fig.˜2(a). All other non-tree edges are shown in light gray. We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in fig.˜2(b). Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in fig.˜2(b). This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$. We refer to such an edge as a non-MST edge, as opposed to an MST edge. By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in fig.˜2(c). For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$. We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in fig.˜2(d).
@@ 4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
$H^{*}_{6}$ in fig.˜4(a) is a hypergraph with similar structure to $H_{6}$ in fig.˜2(a). $H^{*}_{6}$ differs from $H_{6}$ by one vertex $d^{\prime}$ and admits the same set of join trees. Therefore, finding an equivalent graph $G^{\equiv}(H^{*}_{6})$ is sufficient for enumerating the join trees of $H_{6}$. $H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in fig.˜4(b). This means each tree edge in $\hat{T}_{P}$ is already in place, because it cannot slide along the lighter parent. For each non-tree edge, we can identify its destination in constant time using algorithms for finding lowest common ancestors [3] and level ancestors [4], as detailed in algorithm˜3.
@@ 9    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication. lemma˜19 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$. We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$. algorithm˜2 constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
algorithm˜2 starts with enqueuing each edge incident to the root. Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$. The duplication is performed only if $\Delta>0$. lemma˜16 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$. Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
Let $e=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$, $x$ only occurs in the subtree rooted at $r_{u}$. Otherwise, it violates the running intersection property. Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$. By lemma˜17, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold. The tree has monotonic weights when algorithm˜2 terminates. ∎
@@
In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, definition˜7 and definition˜11 guarantee that:
@@ 6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
We consider $\hat{T}_{P}$ in fig.˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in fig.˜4(a). Three of $\hat{T}_{P}$’s non-tree edges are shown in fig.˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed. We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time. Because $\omega(Y,W)=|\{a\}|=1$ is lighter than $\omega(T,W)=\omega(T,U)=|\{a,c\}|=2$, it is removed. $(Y,T)$ is also removed, for it is lighter than its LCA edge $(T,U)$. We compare $(Y,S)$ with $\lambda(Y,S)=\{(P,T),(P,S)\}$. Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$. The intermediate graph with those three non-tree edges processed is shown in fig.˜5(b). The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in fig.˜5(c).
@@
lemma˜22 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges. For example, running algorithm˜1 on the MCS tree $T(H_{6})$ in fig.˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in fig.˜5(c).
@@
One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by lemma˜22. ∎
@@ 5 The Canonical Join Tree of a Berge-Acyclic Query
By proposition˜26, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1. Every spanning tree is a maximum spanning tree, therefore a join tree.
@@ 5.1 Existence & Uniqueness of the Canonical Join Tree
Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], lemma˜31 implies the following corollary.
@@
A spanning tree of $L$ is its join tree by corollary˜27. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
@@
We prove $T_{r}$ is acyclic by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$. Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$. $R_{0}=\{r\}$ contains only the root. The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic. Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$. corollary˜32 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$. Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$. Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$. Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a join tree.
The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$. The canonical tree is unique by corollary˜32. ∎
@@ 5.2 Construction of the Canonical Join Tree
To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$. An empty graph is trivially a subtree of any graph. When the root $r$ is labeled as $r_{1}$, this graph of a single vertex is also trivially a subtree of $T_{r}$. Assuming $G_{k}$ is a subtree of $T_{r}$, algorithm˜1 proceeds to label $r_{k+1}$ and assigns its parent $p(r_{k+1})=r_{p\leq k}$. Suppose for the sake of contradiction that $r_{k+1}$ has a different parent $r_{q}$ in $T_{r}$. Let us consider the possible relationships between $r_{p}$ and $r_{q}$. First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating theorem˜33. Suppose $r_{q}$ is an ancestor of $r_{p}$. By the running intersection property, $r_{q}\cap r_{k+1}$ must contain $r_{p}\cap r_{k+1}$, but that would imply every vertex in $r_{p}\cap r_{k+1}$ is already marked by $r_{q}$ and the algorithm would not have assigned $r_{p}$ as the parent of $r_{k+1}$. Therefore, $r_{p}$ and $r_{q}$ are not descendants of each other. Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$. Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in fig.˜5 of appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in fig.˜5. By lemma˜34, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct. Consider the following possible cases:

### v2

@@ 2 Related Work
Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [28, 34, 30], cost-based pruning from the top down [11, 15], greedy heuristics [5, 14, 38], and randomized search [36]. Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [21, 23], while others avoid Cartesian products [28, 29]. Our algorithms restrict the query plans to those running in linear time for acyclic queries. In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^1] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other. We later define this as the line graph of the query hypergraph in Section 3. our algorithms find maximum spanning trees of the weighted join graph. Several algorithms for ordering binary joins are based on dynamic programming and tabulate shared structures among different plans [28, 29]. This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner. Our enumeration algorithm avoids redundant work by enumerating join trees by edits, i.e., it outputs the difference between consecutive join trees instead of the trees themselves. Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
@@ 13 $R\leftarrow R\setminus\{r_{i}\}$
Figure˜2(a) shows an example hypergraph of size 14. The line graph $L_{6}=L(H_{6})$ is shown in Figure˜2(b). For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph. The edge weight is given by $\omega(Y,U)=|\{a,d\}|=2$. Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$. Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge. The size of the line graph is $|L_{6}|=19$.
@@
The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree. It generates the join tree $T^{G}$ as shown in Figure˜2(c). Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic. The detailed definitions are given by Appendix˜A in Appendix˜A.
Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40]. Algorithm˜1 shows a simplified pseudocode of MCS. We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in Figure˜2(d) with Appendix˜A in Appendix˜A.
@@ 4.1 Equivalent Graph
Figure 2: $T_{W}$ is an MST of the line graph in Figure˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges. $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
@@
Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in Section˜4.1. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
@@
Given the 6-clique line graph $L_{6}$ in Figure˜2(b), we find a rooted MST $T_{W}$ as shown in Figure˜2(a). Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$. We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in Figure˜2(b). Non-tree edges can slide similarly. We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in Figure˜2(a). All other non-tree edges are shown in light gray. We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in Figure˜2(b). Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in Figure˜2(b). This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$. We refer to such an edge as a non-MST edge, as opposed to an MST edge. By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in Figure˜2(c). For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$. We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in Figure˜2(d).
@@ 4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
$H^{*}_{6}$ in Figure˜4(a) is a hypergraph with similar structure to $H_{6}$ in Figure˜2(a). $H^{*}_{6}$ differs from $H_{6}$ by one vertex $d^{\prime}$ and admits the same set of join trees. Therefore, finding an equivalent graph $G^{\equiv}(H^{*}_{6})$ is sufficient for enumerating the join trees of $H_{6}$. $H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in Figure˜4(b). This means each tree edge in $\hat{T}_{P}$ is already in place, because it cannot slide along the lighter parent. For each non-tree edge, we can identify its destination in constant time using algorithms for finding lowest common ancestors [3] and level ancestors [4], as detailed in Algorithm˜3.
@@ 9    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication. Section˜4.2 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$. We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$. Algorithm˜2 constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
Algorithm˜2 starts with enqueuing each edge incident to the root. Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$. The duplication is performed only if $\Delta>0$. Section˜4.2 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$. Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
Let $e=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$, $x$ only occurs in the subtree rooted at $r_{u}$. Otherwise, it violates the running intersection property. Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$. By Section˜4.2, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold. The tree has monotonic weights when Algorithm˜2 terminates. ∎
@@
In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, Section˜4.1 and Section˜4.2 guarantee that:
@@ 6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
We consider $\hat{T}_{P}$ in Figure˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in Figure˜4(a). Three of $\hat{T}_{P}$’s non-tree edges are shown in Figure˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed. We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time. Because $\omega(Y,W)=|\{a\}|=1$ is lighter than $\omega(T,W)=\omega(T,U)=|\{a,c\}|=2$, it is removed. $(Y,T)$ is also removed, for it is lighter than its LCA edge $(T,U)$. We compare $(Y,S)$ with $\lambda(Y,S)=\{(P,T),(P,S)\}$. Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$. The intermediate graph with those three non-tree edges processed is shown in Figure˜5(b). The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in Figure˜5(c).
@@
Section˜4.2 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges. For example, running Algorithm˜1 on the MCS tree $T(H_{6})$ in Figure˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in Figure˜5(c).
@@
One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in Algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by Section˜4.2. ∎
@@ 5 The Canonical Join Tree of a Berge-Acyclic Query
By Section˜5, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1. Every spanning tree is a maximum spanning tree, therefore a join tree.
@@ 5.1 Existence & Uniqueness of the Canonical Join Tree
Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], Section˜5.1 implies the following corollary.
@@
A spanning tree of $L$ is its join tree by Section˜5. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
@@
We prove $T_{r}$ is acyclic by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$. Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$. $R_{0}=\{r\}$ contains only the root. The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic. Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$. Section˜5.1 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$. Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$. Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$. Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a join tree.
The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$. The canonical tree is unique by Section˜5.1. ∎
@@ 5.2 Construction of the Canonical Join Tree
To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$. An empty graph is trivially a subtree of any graph. When the root $r$ is labeled as $r_{1}$, this graph of a single vertex is also trivially a subtree of $T_{r}$. Assuming $G_{k}$ is a subtree of $T_{r}$, Algorithm˜1 proceeds to label $r_{k+1}$ and assigns its parent $p(r_{k+1})=r_{p\leq k}$. Suppose for the sake of contradiction that $r_{k+1}$ has a different parent $r_{q}$ in $T_{r}$. Let us consider the possible relationships between $r_{p}$ and $r_{q}$. First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating Theorem˜33. Suppose $r_{q}$ is an ancestor of $r_{p}$. By the running intersection property, $r_{q}\cap r_{k+1}$ must contain $r_{p}\cap r_{k+1}$, but that would imply every vertex in $r_{p}\cap r_{k+1}$ is already marked by $r_{q}$ and the algorithm would not have assigned $r_{p}$ as the parent of $r_{k+1}$. Therefore, $r_{p}$ and $r_{q}$ are not descendants of each other. Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$. Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in Figure˜5 of Appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in Figure˜5. By Section˜5.2, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct. Consider the following possible cases: